# Augmented Object Intelligence: Making the Analog World Interactable with XR-Objects

### v1

@@ Abstract.
mixed reality, extended reality, augmented reality, augmented objects, spatial computing, object interactions, user interfaces, context menus
@@
Figure 1. XR-Objects allows users to (a) select and interact with real-world objects in AR as if they were digital objects. (b) It uses automatically generated object-based context menus in AR, and objects provide information about themselves, such as a food product. (c, d) The user asks a question about the boiling time, and (e) uses the result to interact with other objects, such as (f) setting up a spatial timer on the pot with the right cooking time. (g) The timer widget is automatically placed on the object in 3D space.
@@ 1. Introduction
The wide availability of machine learning (ML) and computer vision technologies has also led to features that enhance digital interaction with the physical world at our fingertips. Tools like image-based search in Google Lens [^1] Google Lens: https://lens.google and utility-focused Augmented Reality (AR) features in smartphones, such as text copy & paste and real-time translation, are becoming increasingly common. Together, these tools form building blocks that could bring us a little bit closer to a future in which a total understanding of the world and its objects can be applied to our everyday interactions in XR.
In this paper, we explore a novel interaction paradigm we term Augmented Object Intelligence (AOI) that would allow any real object identified by the XR system to reveal digital data associated with the analog object and perform context-appropriate digital actions in a meaningful way. Our system XR-Objects  embodies this idea and aims to demonstrate and investigate ”semantic equality” between real and virtual objects.
@@ 2. Related Work
This section provides an overview of previous work in the realms of user interface (UI) design principles, the interaction with real-world objects, and advancements in AR technology (Speicher et al., 2019). These areas form the foundation upon which our research builds, which aims to enhance the interaction between users and the physical world through XR-Objects.
@@ 2.1. Fundamentals of UI
The challenge of bridging the cognitive gap between human users and computational systems has been a central theme in HCI. Traditional approaches have employed various layers of representations to mediate this interaction, manifesting most recognizably in the UIs of devices ranging from PCs and smartphones to IoT devices and automotive systems (Tidwell et al., 2020). Despite these advancements, the resurgence of command-line interface (CLIs) in contemporary AI interactions, as seen in the usage of prompts with large language models (LLMs) (Suh et al., 2024), suggests a potential oversimplification of user interaction paradigms. This regression shows the necessity of reevaluating our approach to UI design in the age of AI and spatial computing (Suzuki et al., 2023), where the intricacy of human intentions and the computational interpretation thereof demand a more nuanced form of representation. Figure 2 demonstrates that the multimodal AI assistant clearly has the capacity to produce reasonable scene understanding when an image and a prompt is provided as input but it fails in providing an anchored output that ties back to the original multimodal prompt.
@@ 2.2. Blended Reality
The exploration of blended-reality systems is a rapidly developing field within HCI, marked by recent research that integrates real and virtual environments (Hartmann et al., 2019; Lindlbauer and Wilson, 2018; Lindlbauer et al., 2016). While a comprehensive literature review (Auda et al., 2023) exceeds the scope of this section, we draw attention to key themes that emerged within the broader XR literature.
@@
2.3. Detecting and Interacting with AR Objects
@@ 3. XR-Objects Implementation
Given the current technological constraints of AR headsets, particularly their limited developer access to real-time camera streams, we consciously targeted smartphones for our mobile prototype development. We aim to enable anyone to try out our open-source project on their phone. Modern smartphones use similar types of ARM-based mobile chipsets as AR headsets, yielding comparable performance for real-time computer vision tasks, while providing unrestricted access to their high-resolution cameras. This enables our application to identify objects in the user’s environment and overlay digital information directly onto the physical world through the phone’s display thanks to ARCore and ARKit.

### v2

@@ Abstract.
mixed reality; extended reality; augmented reality; augmented objects; spatial computing; user interfaces; context menus
@@
Figure 1. XR-Objects allows users to (a) select and interact with real-world objects in AR as if they were digital objects. (b) Automatically generated object-based AR context menus allow objects to provide information about themselves, such nutritional facts. (c, d, e) For example, a user asks a question about cooking time, and (f,g) uses the result to set a spatial timer widget anchored to the relevant pot in 3D space.
@@ 1. Introduction
The wide availability of machine learning (ML) and computer vision technologies has also led to features that enhance digital interaction with the physical world at our fingertips. Tools like image-based search in Google Lens [^1] Google Lens: https://lens.google and utility-focused Augmented Reality (AR) features in smartphones, such as text copy & paste and real-time translation, are becoming increasingly common. Together, these tools form building blocks that could bring us closer to a future in which a total understanding of the world and its objects can be applied to our everyday interactions in XR.
In this paper, we explore a novel interaction paradigm we term Augmented Object Intelligence (AOI) that would allow any real object identified by the XR system to reveal digital data associated with the analog object and perform context-appropriate digital actions in a meaningful way. Our system XR-Objects  embodies this idea and aims to demonstrate and investigate “semantic equality” between real and virtual objects.
@@ 2. Related Work
This section provides an overview of previous work in the realms of user interface (UI) design principles, blended realiry interactions, and advancements in AR technology (Speicher et al., 2019). These areas form the foundation upon which our research builds, which aims to enhance the interaction between users and the physical world through XR-Objects.
@@ 2.1. Fundamentals of UI
The challenge of bridging the cognitive gap between human users and computational systems has been a central theme in HCI. Traditional approaches have employed various layers of representations to mediate this interaction, manifesting most recognizably in the UIs of devices ranging from PCs and smartphones to IoT devices and automotive systems (Tidwell et al., 2020). Despite these advancements, the resurgence of command-line interface (CLIs) in contemporary AI interactions, as seen in the usage of prompts with large language models (LLMs) (Suh et al., 2024), suggests a potential oversimplification of user interaction paradigms. This regression shows the necessity of reevaluating our approach to UI design in the age of AI and spatial computing (Suzuki et al., 2023), where the intricacy of human intentions and the computational interpretation thereof demand a more nuanced form of representation. Figure 2 demonstrates that the multimodal AI assistant clearly has the capacity to produce reasonable scene understanding when an image and a prompt is provided as input, but it fails in providing an anchored output that ties back to the original multimodal prompt.
@@ 2.2. Blended Reality
The exploration of blended reality systems is a rapidly developing field within HCI, marked by recent research that integrates real and virtual environments (Hartmann et al., 2019; Lindlbauer and Wilson, 2018; Lindlbauer et al., 2016). While a comprehensive literature review (Auda et al., 2023) exceeds the scope of this section, we draw attention to key themes that emerged within the broader XR literature.
@@
2.3. Interacting with Physical Objects in XR
@@ 3. XR-Objects Implementation
Given the current constraints of AR headsets, particularly their limited developer access to real-time camera streams, we consciously targeted smartphones for our mobile prototype development. We aim to enable anyone to try out our open-source project on their phone. Modern smartphones use similar types of ARM-based mobile chipsets as AR headsets, yielding comparable performance for real-time computer vision tasks, while providing unrestricted access to their high-resolution cameras. This enables our application to identify objects in the user’s environment and overlay digital information directly onto the physical world through the phone’s display thanks to ARCore and ARKit.